I need help in drawing out the detailed mechanism for the nitration of salicylic acid...including the reaction equation showing the formation of the nitronium ion from concetrated sulfuric acid and nitric acid.
1
Expert's answer
2013-02-12T10:09:13-0500
Typical nitration syntheses apply so-called "mixed acid", a mixture of concentrated nitric acid and sulfuric acids. This mixture produces the nitronium ion (NO2+), which is the active species in aromatic nitration.
HO - N+ -O- + H-OSO3H = HO+ - N+ -O- + HSO4-
II & I II
O H & O
O
II
HO+ - N+ - O-& =& H2O + O=N+=O
I
H
Electrophilic aromatic substitution, in this case nitration, occurs primarily at the 5 -position in 2-hydroxybenzoic acid (salicylic acid, because the hydroxyl functionality at C-2 orients an incoming electrophile (NO2+) to the positions that are orto and para to it. This can be seen by writing out the resonance structures for this compound, which will place formal negative charges at C-1, C-3 and C-5. C-1 is already occupied, so substitution can only occur at C-3 and C-5. The effect of the carboxylic acid group is deactivating at C-2 (which is occupied by the OH), C-4 and C-6, which again allows only substitution at C-3 and C-5. I suspect that C-5 is somewhat more favored over C-3 simply because of steric effects. That is to say that the OH group at C-2 may interfere somewhat with the approach of the nitro group. This steric hindrance does not occur at C-5, since C-4 and C-6 both have small H-atoms attached.
C6H4CO(OH)2 + HNO3 + H2SO4 = NO2-C6H4CO(OH)2